OpenStudy (anonymous):

Find the Maclaurin series for f(x) using the definition of a Maclaurin series. [Assume that f has a power series expansion. Do not show that Rn(x) → 0.] f(x) = 9(1 − x)^−2

OpenStudy (perl):

is it ok to use wolfram?

OpenStudy (anonymous):

This is beyond me

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\frac{9}{(1-x)^2}\] i bet that is it!

OpenStudy (perl):

there is a trick you can use if you know that 1 / (1-x) = 1 + x + x^2 + x^3 + ... take derivative of both sides the derivative of 1/(1-x) is 1 / (1-x)^2

OpenStudy (anonymous):

forgetting about the 9 for a moment, this is the derivative of the most famousest power series of all

OpenStudy (anonymous):

I'm confused because I'm not quite sure if I'm supposed to take a derivative of this or integrate it to find the series & the infinite sum/